U.S.A. and Canada each produce two goods, an educational product and a medical product. U.S.A is endowed with 2,000 units of labor, and Canada is endowed with 1,500 units of labor. The unit labor requirements are given as:

U.S.A Canada

Unit Labor Requirement for Educational Product

aA=0.6

aA=0.3

Unit Labor Requirement for Medical Product

aM=0.1

aM=0.3

What is the opportunity cost of producing an additional unit of the educational product in the U.S.A & Canada?
Answer #1

Opportunity cost is the value of next best alternative use of a factor. Here factor is labour.If we use one labour and produce 2. Units of education good. And if we could produce 3 units of medical goods with that same labour,then opportunity cost of producing additional unit of education good is=3/2=1.5 medical goods.

Opportunity cost of USA of producing additional unit of educational good=aA/aM=0.6/0.1=6

Opportunity of Canada of producing additional unit of educational good=aA/aM=0.3/0.3=1